Draw a complete electron pushing mechanism for the following reaction.
HOA AN Hg2*, H₂SO4 (dil) H2O, A HO) OH HO (10 pts) A)
Answer #1

n Hat Hy sou u-cc= cc-n H2O, on би Hydration of Alkyne : at Step 1 :- enol formation u-C с= C-C — и + H+ ou ou n и — c— С С—Date Page No. Step 2: Tautomerization 4 H И - С — С. -CH on n 1 fon 04 Protonate Carbon de +4+ И И CH2 с С C-H - Фи h on + H₂